# A Comparison Between Invariant and Equivariant Classical and Quantum Graph Neural Networks

### v2

@@ 3.2 Graph neural network
Classical GNNs take in a collection of graphs $\{\mathcal{G}_{\alpha}\}$ each with nodes $v_{\alpha}^{(i)}\in\mathcal{V}_{\alpha}$ and edges $e_{\alpha}^{(ij)}\in\mathcal{E}_{\alpha}$ where $\mathcal{G}_{\alpha}=\{\mathcal{V}_{\alpha},\mathcal{E}_{\alpha}\}$. Each node $v_{\alpha}^{(i)}$ has an associated feature vector $h_{\alpha}^{(i)}$, and the entire graph has an associated edge attribute tensor $a_{\alpha}^{(ijr)}$ describing $r$ different relationships between node $v_{\alpha}^{(i)}$ and its neighbors $v_{\alpha}^{(j)}$. Here, we can define $\mathcal{N}(i)$ as the set of neighbors of node $v_{\alpha}^{(i)}$ and take $r=1$, as we will only consider one edge attribute, i.e. $a_{\alpha}^{(ijr)}\to a_{\alpha}^{(ij)}$. The edge attributes are typically formed from the features corresponding to each node and its neighbors.
@@
In the layer structure of a GNN, multilayer perceptions (MLPs) are used to update the node features and edge attributes before aggregating, or mean pooling, the updated node features for each graph to make a final prediction. To simplify notation, we will omit the graph index $\alpha$, lower the node index $i$, and introduce a model layer index $l$. The first MLP is the edge MLP $\phi_{e}$ which, at each layer $l$, collects the features $\mathbf{h}_{i}^{l}$, its neighbors’ features $\mathbf{h}_{j}^{l}$, and the edge attribute $a_{ij}$ corresponding to the pair. Once the new edge matrix $m_{ij}$ is formed, we sum along the neighbor dimension $j$ to create a new node feature $\mathbf{m}_{i}$. This extra feature is then added to the original node features $\mathbf{h}_{i}$ before being input into a second node updating MLP $\phi_{h}$ to form new node features $h_{i}^{l+1}$ Kipf and Welling (2016); Gilmer et al. (2017); Satorras et al. (2021). Therefore, a GNN is defined as
$\mathbf{m}_{ij}$ $=\phi_{e}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},a_{ij}),$ (5a)
$\mathbf{m}_{i}$ $=\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij},$ (5b)
$\mathbf{h}_{i}^{l+1}$ $=\phi_{h}(\mathbf{h}_{i}^{l},\mathbf{m}_{i}).$ (5c)
Here, $h_{i}^{l}\in\mathbb{R}^{k}$ is the $k$-th dimensional embedding of node $v_{i}$ at layer $l$ and $\mathbf{m}_{ij}$ is typically referred to as the message passing function. Once the data is sent through the $P$ graph layers of the GNN, the updated nodes $\mathbf{h}_{i}^{P}$ are aggregated via mean pooling for each graph to form a set of final features $\frac{1}{n_{\alpha}}\sum_{i=1}^{n_{\alpha}}\mathbf{h}_{i}^{P}$. These final features are sent through a fully connected neural network (NN) to output the predictions. Typically, a fixed number of hidden features $k=N_{h}$ is defined for both the edge and node MLPs. The described GNN architecture is pictorially shown in the left panel of Fig. 3.
@@ 3.3 SE(2) equivariant graph neural network
For the classical EGNN, the approach used here was informed by the successful implementation of SE(3), or rotational, translational, and permutational, equivariance on dynamical systems and the QM9 molecular dataset Satorras et al. (2021). It should be noted that GNNs are naturally permutation equivariant, in particular invariant, when averaging over the final node feature outputs of the graph layers Thiede et al. (2020). An SE(2) EGNN takes the same form as a GNN, however, the coordinates are equivariantly updated within each graph layer, i.e. $\mathbf{x}_{i}\to\mathbf{x}_{i}^{l}$ where $\mathbf{x}_{i}=(\phi_{i},y_{i})$ in our case. The new form of the network becomes
$\mathbf{m}_{ij}$ $=\phi_{e}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},a_{ij},|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}|),$ (6a)
$\mathbf{m}_{i}$ $=\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij},$ (6b)
$\mathbf{x}_{i}^{l+1}$ $=\mathbf{x}_{i}^{l}+C\sum_{j\neq i}(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l})\phi_{x}(\mathbf{m}_{ij}),$ (6c)
$\mathbf{h}_{i}^{l+1}$ $=\phi_{h}(\mathbf{h}_{i}^{l},\mathbf{m}_{i}).$ (6d)
Since the coordinates $\mathbf{x}_{i}^{l}$ of each node $v_{i}$ are equivariantly evolving, we also introduce a second invariant embedding $|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}|$ based on the equivariant coordinates into the edge MLP $\phi_{e}$. The coordinates $\mathbf{x}_{i}$ are updated by adding the summed difference between coordinate $\mathbf{x}_{i}$ and its neighbors $\mathbf{x}_{j}$ where the added term is enhanced by a factor of $C$, which we take as $C=\frac{1}{\ln(2n_{\alpha})}$, and a coordinate MLP $\phi_{x}$ which takes as input the message passing function $\mathbf{m}_{ij}$. For a proof on the rotational and translational equivariance of $\mathbf{x}_{i}^{l+1}$, see Appendix A. The described EGNN architecture is pictorially shown in the right panel of Fig. 3.
@@ 3.4 Quantum graph neural network
For a QGNN, the input data, as a collection of graphs $\{\mathcal{G}_{\alpha}\}$, is the same as described above. We fix the number of qubits $n$ to be the number of nodes $n_{\alpha}$ in each graph. For the quantum algorithm, we first form a normalized qubit product state from an embedding MLP $\phi_{|\psi^{0}\rangle}$ which takes in the features $\mathbf{h}_{i}$ of each node $v_{i}$ and reduces each of them to a qubit state $|\phi_{|\psi^{0}\rangle}(\mathbf{h}_{i})\rangle\in\mathbb{C}^{2}$ Mernyei et al. (2022). The initial product state describing the system then becomes $|\psi_{\alpha}^{0}\rangle=\bigotimes_{i=1}^{n}|\phi_{|\psi\rangle}(\mathbf{h}_{i})\rangle\in\mathbb{C}^{2^{n}}$, which we normalize by $\sqrt{\langle\psi_{\alpha}^{0}|\psi_{\alpha}^{0}\rangle}$.
Figure 5: $8\times 8$ matrix representations of the coupling and transverse Hamiltonians defined in Eq. (7).
A fully parameterized Hamiltonian can then be constructed based on the adjacency matrix $a_{ij}$ and node features $\mathbf{h}_{i}$ or trainable interaction $\mathcal{W}_{ij}$ and feature $\mathcal{M}_{i}$ weights Verdon et al. (2019). Here, for the coupling term of the Hamiltonian $H_{C}$, we utilize the edge matrix $a_{ij}$ connected to two coupled Pauli-Z operators, $\sigma_{i}^{z}$ and $\sigma_{j}^{z}$, which act on the Hilbert spaces of qubits $i$ and $j$, respectively. Since we embed the quantum state $|\psi^{0}\rangle$ based on the node features $\mathbf{h}_{i}$, we omit the self-interaction term, which utilizes the chosen features applied to the Pauli-Z operator, $\sigma_{i}^{z}$, acting on the Hilbert space of qubit $i$. We also introduce a transverse term $H_{T}$ to each node in the form of a Pauli-X operator, $\sigma_{i}^{x}$ with a fixed or learnable constant coefficient $\mathcal{Q}_{0}$, which we take to be $\mathcal{Q}_{0}=1$. Note that the Hamiltonian $H$ contains $\mathcal{O}(2^{n}\times 2^{n})$ entries due to the Kronecker products between Pauli operators. To best express the properties of each graph, we take the Hamiltonian of the form
@@
where the $8\times 8$ matrix representations of $H_{C}$ and $H_{T}$ are shown in Fig. 5. We generate the unitary form of the operator by complex exponentiating the Hamiltonian with real learnable coefficients $\gamma_{lq}\in\mathbb{R}^{P\times Q}$, which can be thought of as infinitesimal parameters running over $Q=2$ Hamiltonian terms and $P$ layers of the network. Therefore, the QGNN can be defined as
@@
where $U_{\theta}^{l}=(\mathbf{\theta}^{\prime}-i\mathbb{I})(\mathbf{\theta}^{\prime}+i\mathbb{I})^{-1}$ is a parameterized unitary Cayley transformation in which we force $\mathbf{\theta}^{\prime}=\mathbf{\theta}+\mathbf{\theta}^{\dagger}$ to be self-adjoint, i.e. $\mathbf{\theta}^{\prime}=\mathbf{\theta}^{\prime\dagger}$, with $\mathbf{\theta}\in\mathbb{C}^{2^{n}\times 2^{n}}$ as the trainable parameters. The QGNN evolves the quantum state $|\psi^{0}\rangle$ by applying unitarily transformed ansatz Hamiltonians with $Q$ terms to the state over $P$ layers. The final product state $|\psi^{P}\rangle\in\mathbb{C}^{2^{n}}$ is measured for output which is sent to a classical fully connected NN to make a prediction. The analogy between the quantum unitary interaction function $\phi_{u}$ and classical edge MLP $\phi_{e}$, as well as, between the quantum unitary state update function $\phi_{|\psi\rangle}$ and classical node update function $\phi_{h}$ should be clear. For a reduction of the coupling Hamiltonian $H_{C}$ in Eq. (7), see Appendix B. The described QGNN architecture is pictorially shown in the left panel of Fig. 4.
@@ 3.5 Permutation equivariant quantum graph neural Network
The EQGNN takes the same form as the QGNN, however, we aggregate the final elements of the product state $\frac{1}{2^{n}}\sum_{k=1}^{2^{n}}|\psi^{P}_{k}\rangle$ via mean pooling before sending this complex value to a fully connected NN Nguyen et al. (2022); Mernyei et al. (2022); Skolik et al. (2023). See appendix C for a proof of the quantum product state permutation equivariance over the sum of its elements. The resulting EQGNN architecture is shown in the right panel of Fig. 4.
@@ 4 Results and Analysis
For each model, a range of total parameters were tested, however, the overall comparison test was done using the largest optimal number of total parameters for each network. A feed forward NN was used to reduce each network’s graph layered output to a binary one followed by the softmax activation function to obtain the logits in the classical case and the norm of the complex values to obtain the logits in the quantum case. Each model trained over $20$ epochs with the Adam optimizer consisting of a learning rate of $\eta=10^{-3}$ and a cross-entropy loss function. The classical networks trained with a batch size of $64$ and the quantum networks with a batch size of one due to the limited capabilities of broadcasting unitary operators in the quantum APIs. The best model weights were saved when the validation AUC, or area under the curve of the true positive rate (TPR) as a function of the false positive rate (FPR), was maximized after epoch $15$. The results of the training for the largest optimal total number of parameters $|\Theta|\approx 5100$ can be seen in Fig. 6 with more details listed in Table 1.
@@
Table 1: Metric comparison between the classical and quantum graph models.
Model $|\Theta|$ $N_{h}$ P Train ACC Val ACC Test AUC
GNN 5122 10 5 $74.25\%$ $74.80\%$ $\mathbf{63.36\%}$
EGNN 5252 10 4 $73.66\%$ $74.08\%$ $\mathbf{67.88\%}$
QGNN 5156 8 6 $74.00\%$ $73.28\%$ $\mathbf{61.43\%}$
EQGNN 5140 8 6 $74.42\%$ $72.56\%$ $\mathbf{75.17\%}$
@@
Recall that for each model, we vary the number of hidden features $N_{h}$ in the $P$ graph layers. Since we fix the number of nodes $n_{\alpha}=3$ per jet, the hidden feature number $N_{h}=2^{3}=8$ becomes fixed in the quantum case, and therefore, we also varied the parameters of the encoder $\phi_{|\psi^{0}\rangle}$ and decoder NN in the quantum algorithms.
Based on the AUC scores, the EGNN outperformed both the classical and quantum GNN, however, this algortihm was outperformed by the EQGNN with a $7.29\%$ increase in AUC representing the strength of the EQGNN. Although the GNN outperformed the QGNN in the final parameter test by $1.93\%$, the QGNN performed more consistently and outperformed the GNN in the mid parameter range $|\Theta|\in(1500,4000)$. Through the variation of the number of parameters, the AUC scores were recorded for each case where the number of parameters taken for each point corresponds to $|\Theta|\approx\{500,1200,1600,2800,3500,5100\}$ as shown in the right panel of Fig. 7.

### v3

@@ 3.2 Graph Neural Network
Classical GNNs take in a collection of graphs $\{\mathcal{G}_{\alpha}\}$, each with nodes $v_{\alpha}^{(i)}\in\mathcal{V}_{\alpha}$ and edges $e_{\alpha}^{(ij)}\in\mathcal{E}_{\alpha}$, where each graph $\mathcal{G}_{\alpha}=\{\mathcal{V}_{\alpha},\mathcal{E}_{\alpha}\}$ is the set of its corresponding nodes and edges. Each node $v_{\alpha}^{(i)}$ has an associated feature vector $h_{\alpha}^{(il)}$, and the entire graph has an associated edge attribute tensor $a_{\alpha}^{(ijr)}$ describing $r$ different relationships between node $v_{\alpha}^{(i)}$ and its neighbors $v_{\alpha}^{(j)}$. Here, we can define $\mathcal{N}(i)$ as the set of neighbors of node $v_{\alpha}^{(i)}$ and take $r=1$, as we only consider one edge attribute. In other words, the edge attribute tensor $a_{\alpha}^{(ijr)}\to a_{\alpha}^{(ij)}$ becomes a matrix. The edge attributes are typically formed from the features corresponding to each node and its neighbors.
In the layer structure of a GNN, multilayer perceptions (MLPs) are used to update the node features and edge attributes before aggregating, or mean pooling, the updated node features for each graph to make a final prediction. To simplify notation, we omit the graph index $\alpha$, lower the node index $i$, and introduce a model layer index $l$. The first MLP is the edge MLP $\phi_{e}$, which, at each layer $l$, collects the features $\mathbf{h}_{i}^{l}$, its neighbors’ features $\mathbf{h}_{j}^{l}$, and the edge attribute $a_{ij}$ corresponding to the pair. Once the new edge matrix $m_{ij}$ is formed, we sum along the neighbor dimension $j$ to create a new node feature $\mathbf{m}_{i}$. This extra feature is then added to the original node features $\mathbf{h}_{i}$ before being input into a second node updating MLP $\phi_{h}$ to form new node features $h_{i}^{l+1}$ Kipf and Welling (2016); Gilmer et al. (2017); Satorras et al. (2021). Therefore, a GNN is defined as
$\mathbf{m}_{ij}$ $=\phi_{e}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},a_{ij}),$
$\mathbf{m}_{i}$ $=\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij},$ (5)
$\mathbf{h}_{i}^{l+1}$ $=\phi_{h}(\mathbf{h}_{i}^{l},\mathbf{m}_{i}).$
Here, $h_{i}^{l}\in\mathbb{R}^{k}$ is the $k$th-dimensional embedding of node $v_{i}$ at layer $l$, and $\mathbf{m}_{ij}$ is typically referred to as the message-passing function. Once the data are sent through the $P$ graph layers of the GNN, the updated nodes $\mathbf{h}_{i}^{P}$ are aggregated via mean pooling for each graph to form a set of final features $\frac{1}{n_{\alpha}}\sum_{i=1}^{n_{\alpha}}\mathbf{h}_{i}^{P}$. These final features are sent through a fully connected neural network (NN) to output the predictions. Typically, a fixed number of hidden features $k=N_{h}$ is defined for both the edge and node MLPs. The described GNN architecture is pictorially shown in the left panel in Figure 3.
@@ 3.3 SE(2) Equivariant Graph Neural Network
For the classical EGNN, the approach used here was informed by the successful implementation of SE(3), or rotational, translational, and permutational, equivariance on dynamic systems and the QM9 molecular dataset Satorras et al. (2021). It should be noted that GNNs are naturally permutation equivariant, in particular invariant, when averaging over the final node feature outputs of the graph layers Thiede et al. (2020). An SE(2) EGNN takes the same form as a GNN; however, the coordinates are equivariantly updated within each graph layer, i.e., $\mathbf{x}_{i}\to\mathbf{x}_{i}^{l}$ where $\mathbf{x}_{i}=(\phi_{i},y_{i})$ in our case. The new form of the network becomes
$\mathbf{m}_{ij}$ $=\phi_{e}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},a_{ij},|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}|),$
$\mathbf{m}_{i}$ $=\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij},$ (6)
$\mathbf{x}_{i}^{l+1}$ $=\mathbf{x}_{i}^{l}+C\sum_{j\neq i}(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l})\phi_{x}(\mathbf{m}_{ij}),$
$\mathbf{h}_{i}^{l+1}$ $=\phi_{h}(\mathbf{h}_{i}^{l},\mathbf{m}_{i}).$
Since the coordinates $\mathbf{x}_{i}^{l}$ of each node $v_{i}$ are equivariantly evolving, we also introduce a second invariant embedding $|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}|$ based on the equivariant coordinates into the edge MLP $\phi_{e}$. The coordinates $\mathbf{x}_{i}$ are updated by adding the summed difference between the coordinates of $\mathbf{x}_{i}$ and its neighbors $\mathbf{x}_{j}$. This added term is suppressed by a factor of $C$, which we take to be $C(n_{\alpha})=\frac{1}{\ln(2n_{\alpha})}$. The term is further multiplied by a coordinate MLP $\phi_{x},$ which takes as input the message-passing function $\mathbf{m}_{ij}$ between node $i$ and its neighbors $j$. For a proof of the rotational and translational equivariance of $\mathbf{x}_{i}^{l+1}$, see Appendix A. The described EGNN architecture is pictorially shown in the right panel in Figure 3.
@@ 3.4 Quantum Graph Neural Network
For a QGNN, the input data, as a collection of graphs $\{\mathcal{G}_{\alpha}\}$, are the same as described above. We fix the number of qubits $n$ to be the number of nodes $n_{\alpha}$ in each graph. For the quantum algorithm, we first form a normalized qubit product state from an embedding MLP $\phi_{|\psi^{0}\rangle}$, which takes in the features $\mathbf{h}_{i}$ of each node $v_{i}$ and reduces each of them to a qubit state $|\phi_{|\psi^{0}\rangle}(\mathbf{h}_{i})\rangle\in\mathbb{C}^{2}$ Mernyei et al. (2022). The initial product state describing the system then becomes $|\psi_{\alpha}^{0}\rangle=\bigotimes_{i=1}^{n}|\phi_{|\psi\rangle}(\mathbf{h}_{i})\rangle\in\mathbb{C}^{2^{n}}$, which we normalize by $\sqrt{\langle\psi_{\alpha}^{0}|\psi_{\alpha}^{0}\rangle}$.
A fully parameterized Hamiltonian can then be constructed based on the adjacency matrix $a_{ij}$, or trainable interaction weights $\mathcal{W}_{ij}$, and node features $\mathbf{h}_{i}$, or trainable feature weights $\mathcal{M}_{i}$ Verdon et al. (2019). Here, for the coupling term of the Hamiltonian $H_{C}$, we utilize the edge matrix $a_{ij}$ connected to two coupled Pauli-Z operators, $\sigma_{i}^{z}$ and $\sigma_{j}^{z}$, which act on the Hilbert spaces of qubits $i$ and $j$, respectively. Since we embed the quantum state $|\psi^{0}\rangle$ based on the node features $\mathbf{h}_{i}$, we omit the self-interaction term which utilizes the chosen features applied to the Pauli-Z operator, $\sigma_{i}^{z}$, which acts on the Hilbert space of qubit $i$. We also introduce a transverse term $H_{T}$ to each node in the form of a Pauli-X operator, $\sigma_{i}^{x}$, with a fixed or learnable constant coefficient $\mathcal{Q}_{0}$, which we take to be $\mathcal{Q}_{0}=1$. Note that the Hamiltonian $H$ contains $\mathcal{O}(2^{n}\times 2^{n})$ entries due to the Kronecker products between Pauli operators. To best express the properties of each graph, we take the Hamiltonian of the form
@@
where the $8\times 8$ matrix representations of $H_{C}$ and $H_{T}$ are shown in Figure 4. We generate the unitary form of the operator via the complex exponentiating of the Hamiltonian with real learnable coefficients $\gamma_{lq}\in\mathbb{R}^{P\times Q}$, which can be thought of as infinitesimal parameters running over $Q=2$ Hamiltonian terms and $P$ layers of the network. Therefore, the QGNN can be defined as
@@
where $U_{\theta}^{l}=(\boldsymbol{\theta}^{\prime}-i\mathbb{I})(\boldsymbol{\theta}^{\prime}+i\mathbb{I})^{-1}$ is a parameterized unitary Cayley transformation in which we force $\boldsymbol{\theta}^{\prime}=\boldsymbol{\theta}+\boldsymbol{\theta}^{\dagger}$ to be self-adjoint, i.e., $\boldsymbol{\theta}^{\prime}=\boldsymbol{\theta}^{\prime\dagger}$, with $\boldsymbol{\theta}\in\mathbb{C}^{2^{n}\times 2^{n}}$ as the trainable parameters. The QGNN evolves the quantum state $|\psi^{0}\rangle$ by applying unitarily transformed ansatz Hamiltonians with $Q$ terms to the state over $P$ layers.
Figure 4: The $8\times 8$ matrix representations of the coupling and transverse Hamiltonians defined in Equation (7).
The final product state $|\psi^{P}\rangle\in\mathbb{C}^{2^{n}}$ is measured for output, which is sent to a classical fully connected NN to make a prediction. The analogy between the quantum unitary interaction function $\phi_{u}$ and classical edge MLP $\phi_{e}$, as well as between the quantum unitary state update function $\phi_{|\psi\rangle}$ and classical node update function $\phi_{h}$, should be clear. For a reduction in the coupling Hamiltonian $H_{C}$ in Equation (7), see Appendix B. The described QGNN architecture is pictorially shown in the left panel in Figure 5.
@@ 3.5 Permutation Equivariant Quantum Graph Neural Network
The EQGNN takes the same form as the QGNN; however, we aggregate the final elements of the product state $\frac{1}{2^{n}}\sum_{k=1}^{2^{n}}|\psi^{P}_{k}\rangle$ via mean pooling before sending this complex value to a fully connected NN Nguyen et al. (2022); Mernyei et al. (2022); Skolik et al. (2023). See Appendix C for a proof of the quantum product state permutation equivariance over the sum of its elements. The resulting EQGNN architecture is shown in the right panel in Figure 5.
@@ 4 Results and Analysis
For each model, a range of total parameters was tested; however, the overall comparison test was conducted using the largest optimal number of total parameters for each network. A feed-forward NN was used to reduce each network’s graph layered output to a binary one, followed by the softmax activation function to obtain the logits in the classical case and the norm of the complex values to obtain the logits in the quantum case. Each model trained over $20$ epochs with the Adam optimizer consisting of a learning rate of $\eta=10^{-3}$ and a cross-entropy loss function. The classical networks were trained with a batch size of $64$ and the quantum networks with a batch size of one due to the limited capabilities of broadcasting unitary operators in the quantum APIs. After epoch $15$, the best model weights were saved when the validation AUC of the true positive rate (TPR) as a function of the false positive rate (FPR) was maximized. The results of the training for the largest optimal total number of parameters $|\Theta|\approx 5100$ are shown in Figure 6, with more details listed in Table 1.
{adjustwidth}
[…]
@@
[b] Model $|\Theta|$ $N_{h}$ P Train ACC Val ACC Test AUC GNN 5122 10 5 $74.25\%$ $74.80\%$ $\mathbf{63.36\%}$ EGNN 5252 10 4 $73.66\%$ $74.08\%$ $\mathbf{67.88\%}$ QGNN 5156 8 6 $74.00\%$ $73.28\%$ $\mathbf{61.43\%}$ EQGNN 5140 8 6 $74.42\%$ $72.56\%$ $\mathbf{75.17\%}$ 1 The pink color represents the GNN results. The red color represents the EGNN results. The blue color represents the QGNN results. The cyan color represents the EQGNN results. This representation extends to Figure  7.
Table 1: Metric comparison between the classical and quantum graph models 1.
Recall that for each model, we varied the number of hidden features $N_{h}$ in the $P$ graph layers. Since we fixed the number of nodes $n_{\alpha}=3$ per jet, the hidden feature number $N_{h}=2^{3}=8$ was fixed in the quantum case, and, therefore, we also varied the parameters of the encoder $\phi_{|\psi^{0}\rangle}$ and decoder NN in the quantum algorithms.
Based on the AUC scores, the EGNN outperformed both the classical and quantum GNN; however, this algorithm was outperformed by EQGNN with a $7.29\%$ increase in AUC, representing the strength of the EQGNN. Although the GNN outperformed the QGNN in the final parameter test by $1.93\%$, the QGNN performed more consistently and outperformed the GNN in the mid-parameter range $|\Theta|\in(1500,4000)$. Through the variation in the number of parameters, the AUC scores were recorded for each case, where the number of parameters taken for each point corresponded to $|\Theta|\approx\{500,1200,1600,2800,3500,5100\}$, as shown in the right panel in Figure 7.